OpenStudy (anonymous):

Find the inverse function of f(x)= (1+3x)/(5-2x)

OpenStudy (anonymous):

Start with x=(1+3y)/(5-2y) and solve for y.

OpenStudy (anonymous):

This^

OpenStudy (anonymous):

x=(1+3y)/(5-2y) x(5-2y)=(1+3y) 5x-2xy-1=3y y= (5x-2xy-1)/3

OpenStudy (anonymous):

You still have a y on the right side of the =. You need to get it isolated on one side. Try getting all the groups of y on one side, then factoring it out.

OpenStudy (anonymous):

please help me this is where i struggle

OpenStudy (anonymous):

OK, this step is still fine: 5x-2xy-1=3y Get all the groups of y on one side, and the rest of the junk on the other: -2xy-3y=1-5x

OpenStudy (anonymous):

Now factor out y from the left side.

OpenStudy (anonymous):

ok y(-2x-3)=1-5x

OpenStudy (anonymous):

y=1-5x/(-2x-3)

OpenStudy (anonymous):

Yep, and if you wanna get all fancy, you can factor out a -1 from the denominator to get \[f^{-1}(x)=-\frac{1-5x}{2x+3}\]

OpenStudy (anonymous):

Or \[f^{-1}(x) = \frac{5x-1}{2x+3}\]Looks better. Just saying.

OpenStudy (anonymous):

wait so how do u get the inverse?

OpenStudy (anonymous):

Just replace the y and x to each other then solve for y. http://www.mathsisfun.com/sets/function-inverse.html

OpenStudy (anonymous):

@micahwood50 true, true.

OpenStudy (raden):

anyway use this formula will work too : if given f(x) = (ax+b)/(cx+d), so f^(-1)(x) = (-dx+b)/(cx-a) for ilustration,, look here